OpenStudy (anonymous):

Use the method of substitution to solve the linear equation 6a=7-b 6a=5-b

OpenStudy (anonymous):

Solve for either a or b in the first equation, and plug that into the 2nd one!

OpenStudy (anonymous):

still lost

OpenStudy (anonymous):

Something must be wrong with the equation. it can't be right when you add "b" on both side of the equation and have 6a+b=7 6a+b=5 please correct me if I'm wrong

OpenStudy (anonymous):

there is supposed to be an ordered pair,infinitly many solutions or no solution

OpenStudy (anonymous):

I tried several different things and i dont think any are right

OpenStudy (anonymous):

It works

OpenStudy (anonymous):

So solve for a in the first problem (a = 7-b/6)

OpenStudy (anonymous):

then plug your new a into the second problem